OpenStudy (anonymous):

What is the minimum number of 73-ohm resistors that must be connected in parallel to produce an equivalent resistance of 15-ohm or less? Answer is in resistors.

OpenStudy (anonymous):

At least you must have 5 resistors.

OpenStudy (anonymous):

For equal-resistance parallel resistor, you can find the total resistance by dividing R by number of resistor Rt=R/N so N=Rt/R=73/15=4,9.

OpenStudy (ghazi):

\[\frac{ 1 }{ Req }=\frac{ 1 }{ R1 }+\frac{ 1 }{ R2 }+.....\] now R1=R2=....= 73 ohms and Requivalent = 15 \[\frac{ 1 }{ 15 }= \frac{ 1 }{ 73 }+\frac{ 1 }{ 73 }+.....\] which is equal to \[\frac{ 1 }{ 15 }= \frac{ (1+1+1+1+ ....) }{ 73 }=\frac{ n*1 }{ 73 }\] therefore \[n= \frac{ 73 }{15 }=4.8667\] approximately 5